OpenStudy (anonymous):

Solve the mathematical puzzle. S is a three-digit number. Determine the digits of S from the clues. All the digits are odd. The last two digits add to make ten. The first and last digits add to make eight. The first two digits add to make twelve. S is ?

rishavraj (rishavraj):

S=xyz y+z=10 x+z=8 x+y=12 proceed

OpenStudy (anonymous):

I just did the entire process on a piece of notebook paper... The answer is 573, right?

rishavraj (rishavraj):

yuppp

OpenStudy (anonymous):

Thanks!

OpenStudy (tkhunny):

Start enumerating. I'll invent the notation: Digits {123456789}{0123456789}{0123456789} All odd {13579}{13579}{13579} Last 2 make 10: {13579}{{19}{37}{91}{73}} First and Last make 8: Initially: {1357}{{37}{91}{73}} Then: {137}{573}{791} First 2 are 12: {573} Done. Just get ALL the information from each clue.